Solve for the variable. Identify any extraneous solutions. If no solution exists, explain why.
Solve for y: Solve for y: x=y=3 x=> -y+2 y-3
Answer #1

solution. - x = 4+2 sowing for y x(4-3) = 4+2 sy- 3 x = 4+2 xy-y=3x+2 y (x-2) = 3 x +2 y = 3x+2 X -d